OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1, 5) m = -1

OpenStudy (anonymous):

Substitute the x and y coordinates in the equation to find b.

OpenStudy (anonymous):

5 = -1(-1) + b 5 = 1 + b 4 = b y=-x + 4